OpenStudy (anonymous):

@Michele_Laino A box has 8 beads of the same size, but all are different colors. Tamy draws a bead randomly from the box, notes its color, and then puts the bead back in the box. She repeats this 3 times. What is the probability that Tamy would pick a red bead on the first draw, then a white bead, and finally a red bead again? 1/4096 1/512 1/64 1/8

OpenStudy (michele_laino):

Hint: the requested probabilitiy is the product of the single three probabilities

OpenStudy (anonymous):

\[\left(\frac{1}{8}\right)^3=\frac{1}{512} \]

OpenStudy (anonymous):

ohhhh

OpenStudy (michele_laino):

since each of the three probability is equal to 1/8